OpenStudy (anonymous):

Given a hypotenuse of 7 square root 3 in an isosceles right triangle, find the other 2 legs, area, and angle measures. Thanks!

OpenStudy (cwrw238):



OpenStudy (anonymous):

can i use Pythagorean Theorem and do a squared + b squared= 7 square root 3 squared? i got 3 in the end for the other two legs but then i am confused about what type of right triangle it is...it doesn't match a 30-60-90 format or a 45-45-90 format

OpenStudy (cwrw238):

use pythagoras to find x area = 1/2 x^2

OpenStudy (cwrw238):

as its isosceles AND right angled the 2 acute angles are 45 degrees

OpenStudy (cwrw238):

2x^2 = (7 sqrt3)^2

OpenStudy (anonymous):

but the hypotnuse in a 45-45-90 is s square root 2 not 3...

OpenStudy (anonymous):

can i use A= .5bh

OpenStudy (cwrw238):

yep:- but the hypotenuse can be any length you like it can be written a sqrt2 a can be any positive number

OpenStudy (anonymous):

so for length of 2 other legs i got 3 and for area i got 18 does that seem about right?

OpenStudy (cwrw238):

2x^2 = 49*3 = 147 x^2 = 73.5 x = 8.573 to 3 dec places

OpenStudy (anonymous):